# FreeSplat: Generalizable 3D Gaussian Splatting Towards Free-View Synthesis of Indoor Scenes

### v2

@@ Abstract
Empowering 3D Gaussian Splatting with generalization ability is appealing. However, existing generalizable 3D Gaussian Splatting methods are largely confined to narrow-range interpolation between stereo images due to their heavy backbones, thus lacking the ability to accurately localize 3D Gaussian and support free-view synthesis across wide view range. In this paper, we present a novel framework FreeSplat that is capable of reconstructing geometrically consistent 3D scenes from long sequence input towards free-view synthesis. Specifically, we firstly introduce Low-cost Cross-View Aggregation achieved by constructing adaptive cost volumes among nearby views and aggregating features using a multi-scale structure. Subsequently, we present the Pixel-wise Triplet Fusion to eliminate redundancy of 3D Gaussians in overlapping view regions and to aggregate features observed across multiple views. Additionally, we propose a simple but effective free-view training strategy that ensures robust view synthesis across broader view range regardless of the number of views. Our empirical results demonstrate state-of-the-art novel view synthesis peformances in both novel view rendered color maps quality and depth maps accuracy across different numbers of input views. We also show that FreeSplat performs inference more efficiently and can effectively reduce redundant Gaussians, offering the possibility of feed-forward large scene reconstruction without depth priors. Our code will be made open-source upon paper acceptance.
@@ 3 Preliminary
Generalizable 3DGS. Unlike vanilla 3DGS that optimizes per-scene Gaussian primitives, recent generalizable 3DGS [1, 17] predict pixel-aligned Gaussian primitives $\{\bm{\Sigma},\bm{\alpha},\bm{\mathrm{s}}\}$ and depths $\bm{d}$, such that the pixel-aligned Gaussian primitives can be unprojected to 3D coordinates $\bm{\mu}$. The Gaussian parameters are predicted by 2D encoders, which are optimized by the photometric loss through rendering from novel views. However, existing methods are still limited to view interpolation within narrow view range, which leads to inaccurately localized 3D Gaussians that fail to support large scene reconstruction and view extrapolation (cf. Figure 1, 3). To this end, we propose FreeSplat towards global 3D Gaussians reconstruction with accurate localization that supports free-view synthesis.
@@ 4.2 Low-cost Cross-View Aggregation
Efficient 2D Feature Extraction. Given a sparse sequence of posed images $\{\bm{I}^{t}\}_{t=1}^{T}$, we first feed them into a shared 2D backbone to extract multi-scale embeddings $\bm{F}_{e}^{t}$. To decrease and computational cost in plane sweeping, we train a matching net [41] to embed the low-dimensional matching feature $\bm{F}_{m}^{t}$. Unlike [1, 2] which rely on patch-wise transformer-based backbones [42, 43] that can lead to quadratically expensive computations, we leverage pure CNN-based backbones [44, 41] for 2D feature extraction for efficient performance on higher resolution inputs.
@@
Multi-Scale Feature Aggregation. The embedding of the cost volume plays a significant part to accurately localize the 3D Gaussians (cf. Table 5, 6). To this end, inspired by previous depth estimation methods [47, 33], we design an multi-scale encoder-decoder structure, such that to fuse multi-scale image features with the cost volume and propagate the cost volume information to broader receptive fields. Specifically, the multi-scale encoder takes in $\bm{F}_{\mathrm{cv}}^{t}$ and the output is concatenated with $\{\bm{F}_{s}^{t}\}$ before sending into a UNet++ [48]-like decoder to upsample to full resolution and predict a depth candidates map $\bm{D}_{c}^{t}\in\mathbb{R}^{K\times H\times W}$, and Gaussian triplet map $\bm{F}^{t}_{l}\in\mathbb{R}^{C\times H\times W}$. We then predict the depth map through soft-argmax to bound the depth prediction between near and far:
@@ 4.3 Pixel-wise Triplet Fusion
Pixel-wise Alignment. Given the Gaussian triplets $\{\bm{\mu}^{t}_{l},\bm{f}_{l}^{t}\}_{t=1}^{T}$, we start from $t=1$ where the global Gaussians latent is empty. In $t$-th step, we first project the global Gaussian triplet centers $\bm{\mu}_{g}^{t-1}\in\mathbb{R}^{3\times M}$ onto the $t$-th view:
@@
Gaussian Triplet Fusion. After the pixel-wise alignment, we remove the redundant 3D Gaussians through merging the validly aligned triplet pairs. Given a pair $(i,m_{i})\in\rm{\mathcal{F}}^{t}$, to restrict the 3D Gaussian centers to lie between the triplet pair, we weighted sum their center coordinates and sum their weights:
$\bm{\mu}^{t}_{g}(m_{i})=\frac{\bm{\omega}^{t}_{l}(i)\bm{\mu}^{t}_{l}(i)+\bm{\omega}^{t-1}_{g}(m_{i})\bm{\mu}^{t-1}_{g}(m_{i})}{\bm{\omega}^{t}_{l}(i)+\bm{\omega}^{t}_{g}(m_{i})},$ (10)
$\bm{\omega}^{t}_{g}(m_{i})=\bm{\omega}^{t}_{l}(i)+\bm{\omega}^{t-1}_{g}(m_{i}).$ (11)
To aggregate the local and global Gaussian latent features, we update the global Gaussian latent feature through a lightweight GRU network:
$\bm{f}_{g}^{t}(m_{i})=\mathrm{GRU}(\bm{f}_{l}^{t}(i),\bm{f}_{g}^{t-1}(m_{i}))$ (12)
Then the other local Gaussian triplets are appended to the global ones.
@@
$\bm{\Sigma},\bm{\alpha},\bm{\mathrm{s}}=\mathrm{MLP}_{d}(\bm{f}_{g}^{T})$ (13)
and the Gaussian centers $\bm{\mu}=\bm{\mu}_{g}^{T}$. Our proposed fusion method can incrementally integrate the Gaussians with geometrical constraints and learnable GRU network for feature update, which is capable of significantly removing redundant Gaussians and perform post feature aggregation across multiple views, and can be trained with the other framework components end-to-end with eligible computation overhead.
@@ 5.1 Experimental Settings
Implementation Details. Our FreeSplat is trained end-to-end using Adam [52] optimizer with an initial learning rate of $41e-$ 4 for 300k iterations following [1, 2]. Due to the large GPU requirements of [1, 2] given high-resolution images, all input images are resized to $512\times 384$ and batch size is set to 1, to form a fair comparison between different methods. We mainly compare with previous generalizable 3DGS methods in 2, 3, 10 reference view settings to evaluate the models’ performance under different length of input views. For 10 views setting, we also choose target views that are beyond the given sequence of reference views to evaluate the view extrapolation results. Furthermore, we notice that in MVSplat there is a misalignment between the depth plane candidates and the unprojection (more discussion in Appendix 5), which we modified and report their improved version in the following unless otherwise specified.
Table 1: Generalizable Novel View Interpolation results on ScanNet [18]. Except for FreeSplat-fv, other methods are all trained on specific number of views to form a complete comparison. Time(s) indicates the total time of encoding input images and rendering one image.
@@
MVSplat [2] 25.05 0.771 0.258 0.105 393 24.24 0.769 0.273 0.192 590
FreeSplat-spec 27.70 0.826 0.222 0.095 292 27.24 0.820 0.232 0.101 384
FreeSplat-fv 27.18 0.818 0.226 0.097 296 26.79 0.813 0.238 0.104 393
@@
MVSplat [2] 1.178 1966 18.87 0.664 0.417 18.72 0.642 0.444
FreeSplat-spec 0.273 1071 23.06 0.755 0.354 22.30 0.722 0.391
FreeSplat-fv 0.307 958 24.91 0.792 0.278 24.05 0.773 0.304
@@
MVSplat [2] 0.526 0.274 0.396 0.394 0.201 0.625 0.423 0.240 0.558
FreeSplat-spec 0.239 0.121 0.833 0.218 0.108 0.882 0.214 0.122 0.850
FreeSplat-fv 0.220 0.117 0.847 0.209 0.099 0.888 0.159 0.092 0.917
@@ 5.2 Results on ScanNet
View Interpolation Results. On ScanNet, we evaluate the generalizable novel view interpolation results given 2 and 3 reference views as shown in Table 1. Comparing to pixelSplat and MVSplat, our FreeSplat-spec consistently improves PSNR by over $1.67$ dB on 2-views setting and $1.48$ dB on 3-views setting. Although slightly underperforming on SSIM comparing to NeuRay [9], we show significant improvements on PSNR and LPIPS over NeuRay and 300×300$\times$300 × faster inference speed. Moreover, our FreeSplat-fv consistently offers competitive results given arbitrary number of views.
@@
Long Sequence Results. We further evaluate the long sequence results where we sample reference views with length of 10, and compare both view interpolation and extrapolation results as shown in Table 2. The results reveal that generalizable 3DGS methods underperform when given long sequence input images, which is due to the complicated camera trajectories in ScanNet, and the inaccuracy of 3D Gaussian localization that leads to errors when observed beyond the interpolated views. Our FreeSplat-3views still outperforms pixelSplat and MVSplat by over $1.80$ dB PSNR on view interpolation and $1.60$ dB PSNR on view extrapolation results. Through our proposed FVT that can be easily plugged into our model due to our low requirement on GPU, our FreeSplat-fv significantly outperforms our 3-views version by over $1.85$ dB PSNR. Our PTF module can also reduce the number of Gaussians by $45.5\%-51.3\%$, which significantly contributes to the efficiency of 3D Gaussians when given long sequence inputs. The qualitative results are shown in Figure 3, where the results clearly reveal that FreeSplat-spec outperforms MVSplat and pixelSplat in localizing 3D Gaussian and preserving fine-grained details, and FreeSplat-fv further improves on localizing and fusing multi-view Gaussians.
Novel View Depth Estimation Results. We also investigate the correctness of 3D Gaussian localization of different methods through comparing their depth rendering results. We report the Absolute Difference (Abs. Diff), Relative Difference (Rel. Diff), and threshold tolerance $\delta<1.25$ results from novel views in Table 3. We find that FreeSplat significantly outperforms pixelSplat and MVSplat in predicting accurately localized 3D Gaussians, where FreeSplat-fv reaches $91.7\%$ of $\delta<1.25$, outperforming pixelSplat and MVSplat by over $27.0\%$, enabling accurate unsupervised depth estimation from novel views. The improved depth estimation accuracy of FreeSplat-fv highlights the importance of depth estimation in supporting free-view synthesis across broader view range.
@@
MVSplat [2] 21.51 0.749 0.267 0.295 590 16.06 0.644 0.459 0.203 1966
FreeSplat-spec 27.08 0.848 0.175 0.657 415 19.83 0.735 0.361 0.585 1571
FreeSplat-fv 26.28 0.834 0.193 0.655 429 20.99 0.757 0.322 0.665 1478
@@ 5.4 Ablation Study
We conduct a detailed ablation study as shown in Table 5 and Figure 4. The results indicate that cost volume is essential in accurately localizing 3D Gaussians, and our proposed PTF module can consistently contribute to rendering quality and depth estimation results through its post aggregation of Gaussians. Furthermore, our FVT module excels in long sequence reconstruction results as well as novel view depth rendering results, indicating its significance in the unsupervised learning of depth estimation and reconstructing geometrically sound scenes.
@@ A.2 Additional Implementation Details
We set the number of virtual depth planes $K=64$, matching feature dimension $C_{m}=16$, and $d_{near}=0.5,d_{far}=15.0$ for cost volume formulation, and set $\delta=0.05$ in Eq.(8) for pixel-wise alignment. During our training and testing, the interval between reference views is set to 20 for ScanNet and 10 for Replica. To train the 3-views version of pixelSplat [1] on a single NVIDIA RTX A6000 GPU, we change their ViT patch size from $8\times 8$ to $16\times 16$. During inference on the 10 views setting, the epipolor line sampling in pixelSplat and the cross-view attention and cost volume formulation in MVSplat [2] are performed between nearby views similarly as ours to save GPU requirements and form a fair comparison, where for 3-views verion of pixelSplat, MVSplat and FreeSplat we set $N=2$ to match their training settings, and for the free-view version of FreeSplat we set $N=4$.
Table 6: Ablation on cost volume framework design.
3 views
@@
MVSplat [2] 24.24 0.769 0.273 0.394 0.625
MVSplat [2] w/o Matched Depth Candidates 24.02 0.754 0.276 0.422 0.574
FreeSplat-spec w/o Matching Net (MN) 26.46 0.801 0.249 0.254 0.840
FreeSplat-spec w/o Multi-Scale Decoder (MSD) 26.07 0.799 0.281 0.263 0.825
FreeSplat-spec w/o (MN+MSD) 24.94 0.767 0.306 0.300 0.776
FreeSplat-spec 27.24 0.820 0.232 0.218 0.882
A.3 Additional Ablation Study
Figure 5: Mismatched depth plane candidates and unprojection in MVSplat [2].
Comparison with MVSplat. Since MVSplat [2] shares commons with our method in high-level design of building cost volume for generalizable 3DGS, we explore the key design differences that leads to our performance improvements. As shown in Figure 5, we notice that MVSplat uses the predicted depth as the ray distance instead of the distance to the image plane, which results in mismatching between the cost volume and point unprojection. As shown in Table 6, we evaluate the effectiveness of our matching net and the multi-scale decoder to compare with MVSplat. We also evaluate the influence of the mismatching problem on MVSplat. We use the image features $\bm{F}_{s=1}^{t}$ for cost volume formulation when the matching net is removed. We upsample and concatenate the cost volume with $\bm{F}_{s=0}^{t}$ to predict pixel-aligned Gaussian triplets when the multi-scale decoder is removed. The results indicate that removing the matching net (MN) and the multi-scale decoder (MSD) can result in decreased performance on novel view synthesis and depth estimation. Our performance becomes similar as MVSplat when both MN and MSD are removed, which indicates that our cost volume encoding and decoding is more suitable for aggregating more nearby views and can work more effectively on real-world image sequences with complicated camera poses.
Table 7: Ablation on the correlation between performance and the number of input views.
@@ A.3 Additional Ablation Study
FreeSplat-3$\rightarrow$2views 25.70 0.808 0.240 0.293 0.747
FreeSplat-8$\rightarrow$2views 26.35 0.803 0.242 0.228 0.842
FreeSplat-2views 27.70 0.826 0.222 0.239 0.833
FreeSplat-fv 27.18 0.818 0.226 0.220 0.847
FreeSplat-8$\rightarrow$10views 24.92 0.788 0.276 0.153 0.926
FreeSplat-$\rightarrow$10views 24.91 0.792 0.278 0.159 0.917
Correlation between view number and performance. To evaluate the influence of changing the number of input views during inference on our method, we conduct an ablation study as shown in Table 7. The results indicate that training on one specific number of input views can lead to notiable performance decrease when inference on different number of views. However, our free-view training strategy can mitigate such correlation, such that to fit to arbitrary number of input views. On 10-views setting, our FreeSplat-fv can still perform similarly as FreeSplat-8views, indicating that using varying number of input views does not harm the long sequence reconstruction results.

### v3

@@ Abstract
Empowering 3D Gaussian Splatting with generalization ability is appealing. However, existing generalizable 3D Gaussian Splatting methods are largely confined to narrow-range interpolation between stereo images due to their heavy backbones, thus lacking the ability to accurately localize 3D Gaussian and support free-view synthesis across wide view range. In this paper, we present a novel framework FreeSplat that is capable of reconstructing geometrically consistent 3D scenes from long sequence input towards free-view synthesis. Specifically, we firstly introduce Low-cost Cross-View Aggregation achieved by constructing adaptive cost volumes among nearby views and aggregating features using a multi-scale structure. Subsequently, we present the Pixel-wise Triplet Fusion to eliminate redundancy of 3D Gaussians in overlapping view regions and to aggregate features observed across multiple views. Additionally, we propose a simple but effective free-view training strategy that ensures robust view synthesis across broader view range regardless of the number of views. Our empirical results demonstrate state-of-the-art novel view synthesis peformances in both novel view rendered color maps quality and depth maps accuracy across different numbers of input views. We also show that FreeSplat performs inference more efficiently and can effectively reduce redundant Gaussians, offering the possibility of feed-forward large scene reconstruction without depth priors.
@@ 3 Preliminary
Generalizable 3DGS. Unlike vanilla 3DGS that optimizes per-scene Gaussian primitives, recent generalizable 3DGS [1, 17] predict pixel-aligned Gaussian primitives $\{\bm{\Sigma},\bm{\alpha},\bm{\mathrm{s}}\}$ and depths $\bm{d}$, such that the pixel-aligned Gaussian primitives can be unprojected to 3D coordinates $\bm{\mu}$. The Gaussian parameters are predicted by 2D encoders, which are optimized by the photometric loss through rendering from novel views. However, existing methods are still limited to view interpolation within narrow view range, which leads to inaccurately localized 3D Gaussians that fail to support large scene reconstruction and view extrapolation (cf. Figure 1, 4). To this end, we propose FreeSplat towards global 3D Gaussians reconstruction with accurate localization that supports free-view synthesis.
@@ 4.2 Low-cost Cross-View Aggregation
Efficient 2D Feature Extraction. Given a sparse sequence of posed images $\{\bm{I}^{t}\}_{t=1}^{T}$, we first feed them into a shared 2D backbone to extract multi-scale embeddings $\bm{F}_{e}^{t}$ and matching feature $\bm{F}_{m}^{t}$. Unlike [1, 2] which rely on patch-wise transformer-based backbones [41, 42] that can lead to quadratically expensive computations, we leverage pure CNN-based backbones [43, 44] for 2D feature extraction for efficient performance on higher resolution inputs.
@@
Multi-Scale Feature Aggregation. The embedding of the cost volume plays a significant part to accurately localize the 3D Gaussians (cf. Table 5). To this end, inspired by previous depth estimation methods [47, 33], we design an multi-scale encoder-decoder structure, such that to fuse multi-scale image features with the cost volume and propagate the cost volume information to broader receptive fields. Specifically, the multi-scale encoder takes in $\bm{F}_{\mathrm{cv}}^{t}$ and the output is concatenated with $\{\bm{F}_{s}^{t}\}$ before sending into a UNet++ [48]-like decoder to upsample to full resolution and predict a depth candidates map $\bm{D}_{c}^{t}\in\mathbb{R}^{K\times H\times W}$, and Gaussian triplet map $\bm{F}^{t}_{l}\in\mathbb{R}^{C\times H\times W}$. We then predict the depth map through soft-argmax to bound the depth prediction between near and far:
@@
Figure 3: Visual illustration of PTF. The PTF incrementally projects current global Gaussians to input views and computes their pixel-wise distance with local Gaussians. Nearby local Gaussians are then fused using a lightweight Gate Recurrent Unit (GRU) network [49].
@@ 4.3 Pixel-wise Triplet Fusion
Pixel-wise Alignment. Given the Gaussian triplets $\{\bm{\mu}^{t}_{l},\bm{f}_{l}^{t}\}_{t=1}^{T}$, we start from $t=1$ where the global Gaussians latent is empty. In the $t$-th step, we first project the global Gaussian triplet centers $\bm{\mu}_{g}^{t-1}\in\mathbb{R}^{3\times M}$ onto the $t$-th view:
@@
Gaussian Triplet Fusion. After the pixel-wise alignment, we remove the redundant 3D Gaussians through merging the validly aligned triplet pairs. Given a pair $(i,m_{i})\in\rm{\mathcal{F}}^{t}$, we compute the weighted sum of their center coordinates and sum their weights to restrict the 3D Gaussian centers to lie between the triplet pair:
$\bm{\mu}^{t}_{g}(m_{i})=\frac{\bm{\omega}^{t}_{l}(i)\bm{\mu}^{t}_{l}(i)+\bm{\omega}^{t-1}_{g}(m_{i})\bm{\mu}^{t-1}_{g}(m_{i})}{\bm{\omega}^{t}_{l}(i)+\bm{\omega}^{t}_{g}(m_{i})},\quad\text{where}~{}\bm{\omega}^{t}_{g}(m_{i})=\bm{\omega}^{t}_{l}(i)+\bm{\omega}^{t-1}_{g}(m_{i}).$ (10)
We then aggregate the aligned local and global Gaussian latent features through a lightweight GRU network:
$\bm{f}_{g}^{t}(m_{i})=\operatorname{GRU}(\bm{f}_{l}^{t}(i),\bm{f}_{g}^{t-1}(m_{i})),$ (11)
and then append with the other unaligned local Gaussian triplets.
@@
$\bm{\Sigma},\bm{\alpha},\bm{\mathrm{s}}=\operatorname{MLP}_{d}(\bm{f}_{g}^{T})$ (12)
and Gaussian centers $\bm{\mu}=\bm{\mu}_{g}^{\top}$. Our proposed fusion method can incrementally integrate the Gaussians with geometrical constraints and learnable GRU network for feature update. Consequently, our fusion method is capable of significantly removing redundant Gaussians and perform post feature aggregation across multiple views, and can be trained with the other framework components end-to-end with eligible computation overhead.
@@ 5.1 Experimental Settings
Implementation Details. Our FreeSplat is trained end-to-end using Adam [53] optimizer with an initial learning rate of $41e-$ 4 and cosine decay following [2]. Due to the large GPU requirements of [1, 2] given high-resolution images, all input images are resized to $384\times 512$ and batch size is set to 1, to form a fair comparison between different methods. We mainly compare with previous generalizable 3DGS methods in 2, 3, 10 reference view settings, where the distance between input views is fixed, thus evaluating the models’ performance under different view ranges. For 10 views setting, we also choose target views that are beyond the given sequence of reference views to evaluate the view extrapolation results.
Table 1: Generalizable Novel View Interpolation results on ScanNet [18]. FreeSplat-fv is trained with our FVT strategy, and the other methods are all trained on specific number of views to form a complete comparison. Time(s) indicates the total time of encoding input images and rendering one image.
@@
MVSplat [2] 27.27 0.822 0.221 0.117 393 26.68 0.814 0.235 0.192 590
FreeSplat-spec 28.08 0.837 0.211 0.103 278 27.45 0.829 0.222 0.121 382
FreeSplat-fv 27.67 0.830 0.215 0.104 279 27.34 0.826 0.226 0.122 390
@@
MVSplat [2] 1.178 1966 22.78 0.754 0.335 21.60 0.729 0.365
FreeSplat-3views 0.599 882 25.15 0.800 0.278 23.78 0.774 0.309
FreeSplat-fv 0.596 899 25.90 0.808 0.252 24.64 0.786 0.277
@@
MVSplat [2] 0.192 0.106 0.912 0.164 0.079 0.929 0.142 0.080 0.914
FreeSplat-spec 0.157 0.086 0.919 0.161 0.077 0.930 0.120 0.070 0.945
FreeSplat-fv 0.153 0.085 0.923 0.162 0.077 0.928 0.097 0.059 0.961
@@ 5.2 Results on ScanNet
View Interpolation Results. On ScanNet, we evaluate the generalizable novel view interpolation results given 2 and 3 reference views as shown in Table 1. Comparing to pixelSplat and MVSplat, our FreeSplat-spec consistently improves rendering quality and efficiency on 2-views setting and 3-views setting. Although slightly underperforming on SSIM comparing to NeuRay [9], we show significant improvements on PSNR and LPIPS over NeuRay and 300×300$\times$300 × faster inference speed. Moreover, our FreeSplat-fv consistently offers competitive results given arbitrary number of views, and performs more similarly as FreeSplat-spec when number of input views increases.
@@
Long Sequence Results. As shown in Table 2, we further evaluate the long sequence results where we sample reference views with length of 10, and compare both view interpolation and extrapolation results. The results reveal that generalizable 3DGS methods underperform when given long sequence input images, which is due to the complicated camera trajectories in ScanNet, and the inaccuracy of 3D Gaussian localization that leads to errors when observed from wide view ranges. Our FreeSplat-3views significantly outperforms pixelSplat and MVSplat on view interpolation and view extrapolation results. Through our proposed FVT that can be easily plugged into our model due to our low requirement on GPU, our FreeSplat-fv consistently outperforms our 3-views version. Our PTF module can also reduce the number of Gaussians by around $55.0\%$, which becomes indispensable in long sequence reconstruction due to the pixel-wise unprojection nature of generalizable 3DGS. The qualitative results are shown in Figure 4, which clearly reveal that FreeSplat-spec outperforms MVSplat and pixelSplat in localizing 3D Gaussian and preserving fine-grained details, and FreeSplat-fv further improves on localizing and fusing multi-view Gaussians.
Novel View Depth Estimation Results. We also investigate the correctness of 3D Gaussian localization of different methods through comparing their depth rendering results. We report the Absolute Difference (Abs. Diff), Relative Difference (Rel. Diff), and threshold tolerance $\delta<1.25$ results from novel views in Table 3. We find that FreeSplat consistently outperforms pixelSplat and MVSplat in predicting accurately localized 3D Gaussians, where FreeSplat-fv reaches $94.9\%$ of $\delta<1.25$, enabling accurate unsupervised depth estimation on novel views. The improved depth estimation accuracy of FreeSplat-fv highlights the importance of depth estimation in supporting free-view synthesis across broader view range.
@@
MVSplat [2] 26.16 0.840 0.173 0.670 590 18.66 0.717 0.360 0.565 1966
FreeSplat-spec 26.98 0.848 0.171 0.682 423 21.11 0.762 0.312 0.720 1342
FreeSplat-fv 26.64 0.843 0.184 0.682 421 21.95 0.777 0.290 0.742 1346
@@ 5.4 Ablation Study
We conduct a detailed ablation study as shown in Table 5 and Figure 5. The results indicate that: 1) cost volume is essential in accurately localizing 3D Gaussians; 2) our proposed PTF module can consistently contribute to rendering quality and depth estimation results. The PTF module learns to incrementally fuse multi-view 3D Gaussians and contributes significantly when varying number of input views, and serves as a multi-view localization regularization that helps unsupervised depth estimation; 3) Our FVT module excels in long sequence reconstruction quality as well as novel view depth rendering results, which provides stricter constrains on 3D Gaussian localization and can be seamlessly combined with the PTF module to fit to varying length of input views.
@@ 6 Conclusion
7 Acknowledgement
This work is supported by the Agency for Science, Technology and Research (A*STAR) under its MTC Programmatic Funds (Grant No. M23L7b0021).
@@ A.2 Additional Implementation Details
We set the number of virtual depth planes $K=128$, matching feature dimension $C_{m}=64$, and $d_{near}=0.5,d_{far}=15.0$ for cost volume formulation, and set $\delta=0.05$ in Eq.(8) for pixel-wise alignment. To train the 3-views version of pixelSplat [1] on a single NVIDIA RTX A6000 GPU, we change their ViT patch size from $8\times 8$ to $16\times 16$. During inference on the 10 views setting, the epipolar line sampling in pixelSplat and the cross-view attention in MVSplat [2] are performed between nearby views similarly as ours to save GPU requirements and form a fair comparison. For the free-view version of FreeSplat we set the number of nearby views as $N=4$ for training. For the testing of long sequence explicit reconstruction, cost volumes are formed between nearby 8 views.
A.3 Additional Experiments
Table 6: Comparison on computational cost and whole scene reconstruction (30 input views). We report the required GPU for Train / Test, the Encoding Time, the rendering FPS, and PSNR of novel views. - denotes that we are not able to run pixelSplat inference using 30 input views due to its increasing GPU requirement.
@@ A.3 Additional Experiments
pixelSplat-3views 44.9 /   - - - -
MVSplat-3views 34.8 / 44.0 3.004 39 17.57
FreeSplat-3views 16.9 / 21.0 1.191 57 21.33
FreeSplat-fv w/o PTF 42.2 / 21.0 1.006 39 21.82
FreeSplat-fv 42.2 / 21.0 1.205 72 22.32
Computational Cost. As shown in Table 6, we compare the required GPU memory for training and testing, the encoding time, rendering FPS, and PSNR for whole scene reconstruction. pixelSplat-3views and MVSplat-3views already consume 30 50 GB GPU memory for training due to their quadratically increasing GPU memory requirement with respect to the image resolution / sequence length. Therefore, it becomes infeasible to extend their methods to higher resolution inputs or longer sequence training. In comparison, our low-cost framework design enable us to effectively train on long sequence inputs while requiring lesser GPU memory compared to the 3 views version of existing methods. Furthermore, our proposed PTF module can effectively reduce redundant 3D Gaussians, improving rendering speed from 39 to 72 FPS . This becomes increasingly important when reconstructing larger scenes since generalizable 3DGS normally perform pixel-wise unprojection, which can easily result in redundancy in the overlapping regions.
Table 7: Results on RE10K and ACID with 2 input views. We train our model on RE10K, and report its results on RE10K and ACID. denotes that our model is trained on our previously downloaded 9,266 scenes instead of 11,075 scenes used by baselines.
Method RE10K - 2 Views ACID - 2 Views
@@
pixelSplat 25.89 0.858 0.142 27.64 0.830 0.160
MVSplat 26.39 0.869 0.128 28.15 0.841 0.147
Ours 26.41 0.871 0.132 27.94 0.838 0.157
Table 8: Results on RE10K and ACID with 5 input views. The models are trained on RE10K with 5 input views.
Method RE10K - 5 Views ACID - 5 Views
[…]
pixelSplat 24.78 0.850 0.150 26.84 0.833 0.173
MVSplat 25.38 0.866 0.132 27.81 0.863 0.134
Ours 25.95 0.873 0.128 28.35 0.870 0.130
Figure 6: Qualitative Results on RE10K and ACID. We visualize the 2-Views results on RE10K and 5-Views results on ACID.
Experiments on RE10K and ACID. To further evaluate our model’s generalization ability across diverse domains, we train our model on RE10K using 2-View setting and 5-View setting respectively. The results are shown in Table 7, 8 and Figure 6. Note that for the 5-View setting inference, we sample input views with random intervals between 25 and 45 due to the limited sequence lengths in RE10K and ACID. In the 2-View setting, we perform better than pixelSplat and on par as MVSplat on both datasets. In the 5-View setting, we outperform both baselines by a clear margin. We analyze the main causes of the above results as follows:
In the 2-view comparison experiments with the baselines, the image interval between the given stereo images were set to be large. On average, the interval between image stereo is 66 in RE10K and 74 in ACID, which is much larger than our indoor datasets setting (20 for ScanNet and 10 for Replica). Such large interval can result in minimum view overlap between the image stereo, which means that our cost volume can be much sparser and multi-view information aggregation is weakened. In contrast, MVSplat uses a cross-view attention that aggregates multi-view features through a sliding window which does not leverage camera poses. pixelSplat uses a heavy 2D backbone that can potentially become stronger monocular depth estimator. In our 5-view setting, we outperform both baselines by clear margins. This is partially due to the smaller image interval and larger view overlap between nearby views. As a result, our cost volume can effectively aggregate multi-view information, and our PTF module can perform point-level fusion and remove those redundant 3D Gaussians.
Therefore, our model is not specifically designed for highly sparse view inputs, but it is designed as a low-cost model that can easily take in much longer sequences of higher-resolution inputs, that is suitable for indoor scene reconstruction. Comparing to RE10K and ACID, real-world indoor scene sequences usually contain more complicated camera rotations and translations, which results in the requirement of more dense observations to reconstruct the 3D scenes with high completeness and accurate geometry. Consequently, our model is targeting the fast indoor scene reconstruction with keyframe inputs, which contain long sequences of high-resolution images, while existing works struggle to extend to such setting as evaluated in our main paper.
Table 9: Comparison with SurfelNeRF. We compare with SurfelNeRF on the same sequences as their test set. denotes the rendering speed is reported from SurfelNeRF.
[…]
SurfelNeRF 3.242 5 24.20 0.694 0.477
FreeSplat-fv 0.302 224 27.06 0.818 0.223
Figure 7: Qualitative Comparison with SurfelNeRF.
Comparison with SurfelNeRF. We further compare with SurfelNeRF as shown in Table 9 and Figure 7. We evaluate on the same novel views as theirs, sampling input views along their input sequences with an interval of 20 between nearby views. Note that the number of input views changes when the input length changes, while our FreeSplat-fv can seamlessly conduct inference with arbitrary numbers of inputs. Our method performs significantly better than SurfelNeRF in both rendering quality and efficiency. Our end-to-end framework jointly learns depths and 3DGS using an MVS-based backbone, while SurfelNeRF relies on depths and does not aggregate multi-view features to assist their surfel feature prediction.